# SMOGAN: Synthetic Minority Oversampling with GAN Refinement for Imbalanced Regression

### v1

@@ 4.2 Metrics
We evaluate SMOGAN against SMOGN using multiple metrics that measure performance on both frequent and rare target values.
@@ 4.2.2 Squared error-relevance area (SERA)
SERA is a metric that provides a flexible way to evaluate models under non-uniform domain preferences. Let $\phi(y)\in[0,1]$ be a relevance function that assigns higher scores to rare target values [46]. Then for any relevance threshold $t$, let
@@ 4.2.3 Precision, recall and f-measure for regression
Precision and recall were first adapted to regression to assess how well predictions focus on user-specified rare regions rather than minimizing overall error [54, 47, 10].
@@ 4.3 Experimental Framework
For each dataset, we employed 25 different random splits. Each split partitioned the dataset into 80% for training and 20% for testing. Models were trained on the 80% portion and evaluated on the remaining 20%. For our evaluation metrics, we utilized the SERA implementation from the ImbalancedLearningRegression Python package [59]. The SMOGN method was implemented using the package developed by [37].
We trained a TabNet regressor [3] on each training set, holding out 20% for validation. Training ran for up to 1000 epochs with the Adam optimizer (learning rate = 0.0001). The model checkpoint with the lowest validation RMSE was retained and evaluated on the held-out test set. TabNet has six transformer blocks (one initial block plus five decision blocks). Each block contains four fully connected layers of 64 neurons, five attention layers of 16 neurons each, and finally, a single output layer with 1 neuron. We also trained a Random Forest regressor using scikit-learn’s default implementation on the full training set, and then evaluated it directly on the held-out test split [11].
@@ 5.1 Principal component analysis
Figure 5 presents t-SNE scatter plots of the original minority samples alongside SMOGN-generated and SMOGAN-refined samples, using perplexities of 20, 40 and 60. Unlike PCA, which is a linear projection that maximizes global variance, t-SNE is a nonlinear embedding that converts high-dimensional similarities into probabilities and optimizes a 2D map by minimizing their Kullback–Leibler divergence [56]. Perplexity, roughly the effective number of nearest neighbours, balances local versus broader structure, and the axes “t-SNE Dimension 1” and “t-SNE Dimension 2” are simply the two coordinates of this embedding. As a result, t-SNE highlights cluster separation and local density differences that PCA may obscure.
@@ 5.3 Feature Correlation Analysis
[…]
@@
[…]
@@ 6 Results
We present the results of comparing SMOGAN with the state-of-the-art oversampler, SMOGN and a baseline (no oversampling) across various performance dimensions. Note that the SMOGN dataset comprises the original data plus synthetic samples generated by SMOGN, while the SMOGAN dataset comprises the original data plus refined synthetic samples after DistGAN filtration. In Figure 8, we conduct aggregated pairwise comparisons at relevance thresholds $t_{R}=0.8$ for each evaluation metric (RMSE, SERA, F-measure) to demonstrate how SMOGAN performs relative to SMOGN and the baseline.
[…]
(a) SMOGAN vs. SMOGN (TabNet)
[…]
(b) SMOGAN vs. Baseline (TabNet)
[…]
(c) SMOGN vs. Baseline (TabNet)
[…]
(d) SMOGAN vs. SMOGN (Random Forest)
[…]
(e) SMOGAN vs. Baseline (Random Forest)
[…]
(f) SMOGN vs. Baseline (Random Forest)
Figure 8: Pairwise comparisons for TabNet (top row) and Random Forest (bottom row); in each subplot, bars from top to bottom represent F-measure, SERA, and RMSE, respectively. The lighter color represents the total number of wins for each method, while the darker portion indicates wins that are statistically significant.
@@
For each dataset and each pairwise comparison, a winner is determined based on outcomes across 25 random splits, so each dataset contributes one win per threshold and two wins in total across both thresholds. Specifically, for any two methods under comparison (for example, SMOGAN vs. SMOGN or SMOGAN vs. baseline), the method that prevails in the majority of splits is declared the winner. Next, we perform the Wilcoxon signed-rank test for each pairwise comparison. We compute the differences between the 25-split metric values of the two methods to form a difference vector, and then the test assesses whether these differences are statistically significant at an alpha level of 0.05, corresponding to a 5% risk of incorrectly concluding that a difference exists when there is none [58].
@@
In Figure 8, the lighter color represents the total number of wins for each method, while the darker portion indicates wins that are statistically significant. SMOGAN outperforms both SMOGN and the baseline across all three metrics (RMSE, SERA and F-measure) for both TabNet and Random Forest. It not only records a higher total number of wins in every comparison but also contributes the majority of statistically significant wins, demonstrating that adversarial refinement produces synthetic samples of substantially higher fidelity and SMOGAN’s refined data yield consistently better alignment with the true distribution and thus stronger predictive accuracy.
7 Discussion and Conclusion
Our experiments show that even when arbitrary thresholds are used to define minority samples, SMOGAN enhances generalization and predictive accuracy, particularly for extreme values, by generating high-quality synthetic data. In practice, SMOGAN can be integrated into any data-level oversampling pipeline to improve model performance. SMOGAN’s filtering framework addresses low-quality synthetic data by integrating our DistGAN adversarial filtering layer that aligns generated samples with the true joint feature–target distribution. Qualitative analyses (PCA, t-SNE, correlation alignment, statistical moments) and quantitative benchmarks (RMSE, SERA, F-measure) consistently demonstrate that SMOGAN produces more realistic data and achieves superior predictive performance compared to the base oversampler and the baseline model.
@@ 7 Discussion and Conclusion
Nonetheless, SMOGAN’s adversarial refinement adds complexity and computational overhead and requires fine-tuning of GAN-specific hyperparameters for optimal results. Future work will focus on automated fine-tuning strategies for relevance thresholds, more efficient adversarial training schedules, and extensions to high-dimensional or multimodal datasets. Overall, SMOGAN offers a general, data-driven approach that improves the fidelity and utility of synthetic samples for imbalanced regression tasks.

### v2

@@ 4.2 Metrics
We evaluate our method using multiple metrics that measure performance on both frequent and rare target values.
@@ 4.2.2 Squared error-relevance area (SERA)
SERA is a metric that provides a flexible way to evaluate models under non-uniform domain preferences. Let $\phi(y)\in[0,1]$ be a relevance function that assigns higher scores to rare target values (Ribeiro and Moniz, 2020). Then for any relevance threshold $t$, let
@@ 4.2.3 Precision, recall and f-measure for regression
Precision and recall were first adapted to regression to assess how well predictions focus on user-specified rare regions rather than minimizing overall error (Torgo and Ribeiro, 2009; Ribeiro, 2011; Branco et al., 2019).
@@ 4.3 Experimental Framework
For each dataset, we employed 25 different random splits. Each split partitioned the dataset into 80% for training and 20% for testing. Models were trained on the 80% portion and evaluated on the remaining 20%. For our evaluation metrics, we utilized the SERA implementation from the ImbalancedLearningRegression Python package (Wu et al., 2022). The SMOGN method was implemented using the package developed by (Kunz, 2020).
We trained a TabNet regressor (Arık and Pfister, 2021) on each training set, holding out 20% for validation. Training ran for up to 1000 epochs with the Adam optimizer (learning rate = 0.0001). The model checkpoint with the lowest validation RMSE was retained and evaluated on the held-out test set. TabNet has six transformer blocks (one initial block plus five decision blocks). Each block contains four fully connected layers of 64 neurons, five attention layers of 16 neurons each, and finally, a single output layer with 1 neuron. We also trained a Random Forest regressor using scikit-learn’s default implementation on the full training set, and then evaluated it directly on the held-out test split (Breiman, 2001).
@@ 5.1 Principal component analysis
Figure 5 presents t-SNE scatter plots of the original minority samples alongside SMOGN-generated and SMOGAN-refined samples, using perplexities of 20, 40 and 60. Unlike PCA, which is a linear projection that maximizes global variance, t-SNE is a nonlinear embedding that converts high-dimensional similarities into probabilities and optimizes a 2D map by minimizing their Kullback-Leibler divergence (van der Maaten and Hinton, 2008). Perplexity, roughly the effective number of nearest neighbours, balances local versus broader structure, and the axes “t-SNE Dimension 1” and “t-SNE Dimension 2” are simply the two coordinates of this embedding. As a result, t-SNE highlights cluster separation and local density differences that PCA may obscure.
@@ 5.3 Feature Correlation Analysis
(a) Difference $|\text{Real}-\text{SMOGN}|$
(b) Difference $|\text{Real}-\text{SMOGAN}|$
@@
Figure 6 visualizes the absolute difference heatmaps between the real minority correlation matrix and those of SMOGN (a) and SMOGAN (b). A lower Frobenius norm indicates a closer match to the real data’s correlation structure. In our experiments, the norm for the $|\text{Real}-\text{SMOGN}|$ matrix is 2.5060 versus 0.7130 for $|\text{Real}-\text{SMOGAN}|$, representing a 72% reduction in overall discrepancy. This substantial drop shows that SMOGAN’s adversarial refinement preserves inter-feature relationships far better than the initial SMOGN oversampler.
@@ 5.4 Statistical Property Analysis
Figure 8: SMoGAN performance comparison across methods and metrics for TabNet model. Blue bars represent SMoGAN wins, orange bars represent competitor wins. Numbers show total wins with significant wins in parentheses. Higher values indicate better performance.
Figure 9: SMoGAN performance comparison across methods and metrics for random forest model. Blue bars represent SMoGAN wins, orange bars represent competitor wins. Numbers show total wins with significant wins in parentheses. Higher values indicate better performance.
@@ 6 Results
We present the results of comparing SMOGAN with the state-of-the-art oversamplers, SMOGN, WSMOTER, RO and G-SMOTE, as well as baseline (no oversampling) and GAN-only approach across various performance dimensions. Note that the SMOGN dataset comprises the original data plus synthetic samples generated by SMOGN, while the SMOGAN dataset comprises the original data plus refined synthetic samples after DistGAN filtration. We conduct aggregated pairwise comparisons at relevance thresholds $t_{R}=0.8$ for each evaluation metric (RMSE, SERA, F-measure).
@@
For each dataset and each pairwise comparison, a winner is determined based on outcomes across 25 random splits, so each dataset contributes one win per threshold and two wins in total across both thresholds. Specifically, for any two methods under comparison (for example, SMOGAN vs. SMOGN), the method that prevails in the majority of splits is declared the winner. Next, we perform the Wilcoxon signed-rank test for each pairwise comparison. We compute the differences between the 25-split metric values of the two methods to form a difference vector, and then the test assesses whether these differences are statistically significant at an alpha level of 0.05, corresponding to a 5% risk of incorrectly concluding that a difference exists when there is none (Wilcoxon, 1945).
@@
In Figures 8 and 9, the numbers outside the parentheses indicate the total number of wins for each method, while the numbers inside the parentheses represent statistically significant wins. SMOGAN outperforms all competitors across all three metrics (RMSE, SERA, and F-measure) for both TabNet and Random Forest (except against WSMOTER for SERA in the Random Forest model). It not only records a higher total number of wins in every comparison but also contributes the majority of statistically significant wins, demonstrating that adversarial refinement produces synthetic samples of substantially higher fidelity and SMOGAN’s refined data yield consistently better alignment with the true distribution and thus stronger predictive accuracy.
Two key comparisons highlight SMOGAN’s effectiveness. First, SMOGAN significantly outperforms the GAN-only method, demonstrating that it is not simply the GAN process that generates suitable data. Providing the GAN with SMOGN-generated starting points enables more consistent synthetic data generation compared to purely adversarial approaches. Second, SMOGAN’s superiority over SMOGN alone shows that the adversarial loss function successfully moves the synthetic data toward more realistic distributions. This confirms the evidence presented in Section 5, where SMOGAN data demonstrates higher fidelity reproduction of natural minority patterns compared to traditional oversampling methods.
7 Limitations and Discussion
@@
SMOGAN’s adversarial refinement adds complexity and computational overhead and requires fine-tuning of GAN-specific hyperparameters for optimal results. Future work will focus on automated fine-tuning strategies for relevance thresholds, more efficient adversarial training schedules, and extensions to high-dimensional or multimodal datasets. Overall, SMOGAN offers a general, data-driven approach that improves the fidelity and utility of synthetic samples for imbalanced regression tasks.
8 Conclusion
Our experiments show that even when user-selected thresholds are used to define minority samples, SMOGAN enhances generalization and predictive accuracy, particularly for extreme values, by generating high-quality synthetic data. In practice, SMOGAN can be integrated into any data-level oversampling pipeline to improve model performance. SMOGAN’s filtering framework addresses low-quality synthetic data by integrating our DistGAN adversarial filtering layer that aligns generated samples with the true joint feature-target distribution. Qualitative analyses (PCA, t-SNE, correlation alignment, statistical moments) and quantitative benchmarks (RMSE, SERA, F-measure) consistently demonstrate that SMOGAN produces more realistic data and achieves superior predictive performance compared to the base oversampler and the baseline model.